A minor league baseball team plays 96 games in a season if the team won 16 more than three times as many games as they lost how many wins and losses did the team have?
Answer by fractalier(6550)

Call the number of games they won and lost, x and y. Then we write
x + y = 96 and
x = 3y + 16
we can substitute the second into the first and get
(3y+16) + y = 96
4y + 16 = 96
4y = 80
y = 20
which means they won
x = 76
Not a bad season, really...76-20...